ID |
Date |
Author |
Type |
Category |
Subject |
16409
|
Fri Oct 15 20:53:49 2021 |
Koji | Summary | General | Vent Prep |
From the IFO point of view, all look good and we are ready for venting from Mon Oct 18 9AM |
16408
|
Fri Oct 15 17:17:51 2021 |
Koji | Summary | General | Vent Prep |
I took over the vent prep: I'm going through the list in [ELOG 15649] and [ELOG 15651]. I will also look at [ELOG 15652] at the day of venting.
- IFO alignment: Two arms are already locking. The dark port beam is well overlapped. We will move PRM/SRM etc. So we don't need to worry about them. [Attachment 1]
scripts>z read C1:SUS-BS_PIT_BIAS C1:SUS-BS_YAW_BIAS
-304.7661529521767
-109.23924626857811
scripts>z read C1:SUS-ITMX_PIT_BIAS C1:SUS-ITMX_YAW_BIAS
15.534616817500943
-503.4536332290159
scripts>z read C1:SUS-ITMY_PIT_BIAS C1:SUS-ITMY_YAW_BIAS
653.0100945988496
-478.16260735781225
scripts>z read C1:SUS-ETMX_PIT_BIAS C1:SUS-ETMX_YAW_BIAS
-136.17863332517527
181.09285307121306
scripts>z read C1:SUS-ETMY_PIT_BIAS C1:SUS-ETMY_YAW_BIAS
-196.6200333695437
-85.40819256078339
- IMC alignment: Locking nicely. I ran WFS relief to move the WFS output on to the alignment sliders. All the WFS feedback values are now <10. Here is the slider snapshots. [Attachment 2]
- PMC alignmnet: The PMC looked like it was quite misaligned -> aligned. IMC/PMC locking snapshot [Attachment 3]
Arm transmissions:
scripts>z avg 10 C1:LSC-TRX_OUT C1:LSC-TRY_OUT
C1:LSC-TRX_OUT 0.9825591325759888
C1:LSC-TRY_OUT 0.9488834202289581
- Suspension Status Snapshot [Attachment 4]
- Anchal aligned the OPLEV beams [ELOG 16407]
I also checked the 100 days trend of the OPLEV sum power. The trend of the max values look flat and fine. [Attachment 5] For this purpose, the PRM and SRM was aligned and the SRM oplev was also aligned. The SRM sum was 23580 when aligned and it was just fine (this is not so visible in the trend plot).
- The X and Y green beams were aligned for the cavity TEM00s. Y end green PZT values were nulled. The transmission I could reach was as follows.
>z read C1:ALS-TRX_OUTPUT C1:ALS-TRY_OUTPUT
0.42343354488901286
0.24739624058377277
It seems that these GTRX and GTRY seemed to have crosstalk. When each green shutters were closed the transmissino and the dark offset were measured to be
>z read C1:ALS-TRX_OUTPUT C1:ALS-TRY_OUTPUT
0.41822833190834546
0.025039383697636856
>z read C1:ALS-TRX_OUTPUT C1:ALS-TRY_OUTPUT
0.00021112720155274818
0.2249448773499293
Note that Y green seemed to have significant (~0.1) of 1st order HOM. I don't know why I could not transfer this power into TEM00. I could not find any significant clipping of the TR beams on the PSL table PDs.
- IMC Power reduction
Now we have nice motorized HWP. sitemap -> PSL -> Power control
== Initial condition == [Attachment 6]
C1:IOO-HWP_POS 38.83
Measured input power = 0.99W
C1:IOO-MC_RFPD_DCMON = 5.38
== Power reduction == [Attachment 7]
- The motor was enabled upon rotation on the screen
C1:IOO-HWP_POS 74.23
Measured input power = 98mW
C1:IOO-MC_RFPD_DCMON = 0.537
- Then, the motor was disabled
- Went to the detection table and swapped the 10% reflector with the 98% reflector stored on the same table. [Attachments 8/9]
After the beam alignment the MC REFL PD received about the same amount of the light as before.
C1:IOO-MC_RFPD_DCMON = 5.6
There is no beam delivered to the WFS paths.
CAUTION: IF THE POWER IS INCREASED TO THE NOMINAL WITH THIS CONFIGURATION, MC REFL PD WILL BE DESTROYED.
- The IMC can already be locked with this configuration. But for the MC Autolocker, the MCTRANS threshold for the autolocker needs to be reduced as well.
This was done by swapping a line in /opt/rtcds/caltech/c1/scripts/MC/AutoLockMC.init
# BEFORE
/bin/csh ./AutoLockMC.csh >> $LOGFILE
#/bin/csh ./AutoLockMC_LowPower.csh >> $LOGFILE
--->
# AFTER
#/bin/csh ./AutoLockMC.csh >> $LOGFILE
/bin/csh ./AutoLockMC_LowPower.csh >> $LOGFILE
Confirmed that the autolocker works a few times by toggling the PSL shutter. The PSL shutter was closed upon the completion of the test
- Walked around the lab and checked all the bellows - the jam nuts are all tight, and I couldn't move them with my hands. So this is okay according to the ancient tale by Steve.
|
Attachment 1: Screenshot_2021-10-15_17-36-00.png
|
|
Attachment 2: Screenshot_2021-10-15_17-39-58.png
|
|
Attachment 3: Screenshot_2021-10-15_17-42-20.png
|
|
Attachment 4: Screenshot_2021-10-15_17-46-13.png
|
|
Attachment 5: Screenshot_2021-10-15_18-05-54.png
|
|
Attachment 6: Screen_Shot_2021-10-15_at_19.45.05.png
|
|
Attachment 7: Screen_Shot_2021-10-15_at_19.47.10.png
|
|
16407
|
Fri Oct 15 16:46:27 2021 |
Anchal | Summary | Optical Levers | Vent Prep |
I centered all the optical levers on ITMX, ITMY, ETMX, ETMY, and BS to a position where the single arm lock on both were best aligned. Unfortunately, we are seeing the TRX at 0.78 and TRY at 0.76 at the most aligned positions. It seems less power is getting out of PMC since last month. (Attachment 1).
Then, I tried to lock PRMI with carrier with no luck. But I was able to see flashing of up to 4000 counts in POP_DC. At this position, I centered the PRM optical lever too (Attachment 2). |
Attachment 1: Screen_Shot_2021-10-15_at_4.34.45_PM.png
|
|
Attachment 2: Screen_Shot_2021-10-15_at_4.45.31_PM.png
|
|
Attachment 3: Screen_Shot_2021-10-15_at_4.34.45_PM.png
|
|
Attachment 4: Screen_Shot_2021-10-15_at_4.34.45_PM.png
|
|
16406
|
Fri Oct 15 12:14:27 2021 |
Ian MacMillan | Update | General | Kicking optics in freeSwing measurment |
[Ian, Anchal]
we ran the free swinging test last night and the results match up with in 1/10th of a Hz. We calculated the peak using the getPeakFreqs2 script to find the peaks and they are close to previous values from 2016.
In attachment 1 you will see the results of the test for each optic.
The peak values are as follows:
Optic |
POS (Hz) |
PIT (Hz) |
YAW (Hz) |
SIDE (Hz) |
PRM |
0.94 |
0.96 |
0.99 |
0.99 |
MC2 |
0.97 |
0.75 |
0.82 |
0.99 |
ETMY |
0.98 |
0.98 |
0.95 |
0.95 |
MC1 |
0.97 |
0.68 |
0.80 |
1.00 |
ITMX |
0.95 |
0.68 |
0.68 |
0.98 |
ETMX |
0.96 |
0.73 |
0.85 |
1.00 |
BS |
0.99 |
0.74 |
0.80 |
0.96 |
ITMY |
0.98 |
0.72 |
0.72 |
0.98 |
MC3 |
0.98 |
0.77 |
0.84 |
0.97 |
The results from 2016 can be found at: /cvs/cds/rtcdt/caltech/c1/scripts/SUS/PeakFit/parameters2.m |
Attachment 1: 20211015_Kicktest_plot.pdf
|
|
16405
|
Thu Oct 14 20:16:22 2021 |
Yehonathan | Update | General | PRMI free swinging |
{Yehonathan, Raj}
We aligned the IFO in the PRMI state and let it swing freely. |
16404
|
Thu Oct 14 18:30:23 2021 |
Koji | Summary | VAC | Flange/Cable Stand Configuration |
Flange Configuration for BHD
We will need total 5 new cable stands. So Qty.6 is the number to be ordered.
Looking at the accuglass drawing, the in-vaccum cables are standard dsub 25pin cables only with two standard fixing threads.
https://www.accuglassproducts.com/sites/default/files/PDF/Partpdf/110070_3.pdf
For SOSs, the standard 40m style cable bracket works fine. https://dcc.ligo.org/D010194-x0
However, for the OMCs, we need to make the thread holes available so that we can mate DB25 male cables to these cables.
One possibility is to improvise this cable bracket to suspend the cables using clean Cu wires or something. I think we can deal with this issue in situ.
Ha! The male side has the 4-40 standoff (jack) screws. So we can hold the male side on the bracket using the standoff (jack) screws and plug the female cables. OK! The issue solved!
https://www.accuglassproducts.com/sites/default/files/PDF/Partpdf/110029_3.pdf |
Attachment 1: 40m_flange_layout_20211014.pdf
|
|
16403
|
Thu Oct 14 16:38:26 2021 |
Ian MacMillan | Update | General | Kicking optics in freeSwing measurment |
[Ian, Anchal]
We are going to kick the optics tonight at 2am.
The optics we will kick are the PRM BS ITMX ITMY ETMX ETMY
We will kick each one once and record for 2000 seconds and the log files will be placed in users/ian/20211015_FreeSwingTest/logs. |
16402
|
Thu Oct 14 13:40:49 2021 |
Yehonathan | Summary | SUS | PRM and BS Angular Actuation transfer function magnitude measurements |
Here is a side by side comparison of the PRMI sensing matrix using PRM/BS actuation (attachment 1) and ITMs actuation (attachment 2). The situation looks similar in both cases. That is, good orthogonality on REFL55 and bad seperation in the rest of the RFPDs.
Quote: |
should compare side by side with the ITM PRMI radar plots to see if there is a difference. How do your new plots compare with Gautam's plots of PRMI?
|
|
Attachment 1: BSPRM_Actuation_Radar_plots.png
|
|
Attachment 2: ITM_Actuation_Radar_plots.png
|
|
16401
|
Thu Oct 14 11:25:49 2021 |
Yehonathan | Update | PSL | PMC unlocked |
{Yehonathan, Anchal}
I went to get a sandwich around 10:20 AM and when I came back BS was moving like crazy. We shutdown the watchdog.
We look at the spectra of the OSEMs (attachment 1). Clearly, the UR sensing is bad.
We took the BS sattelite box out. Anchal opened the box and nothing seemed wrong visually. We returned the box and connected it to the fake OSEM box. The sensor spectra seemed normal.
We connected the box to the vacuum chamber and the spectra is still normal (attachment 2).
We turn on the coils and the motion got damped very quickly (RMS <0.5mV).
Either the problem was solved by disconnecting and connecting the cables or it will come back to haunt us.
|
Attachment 1: BS_OSEM_Sensor_PSD.pdf
|
|
Attachment 2: BS_OSEM_Sensor_PSD_AfterReconnectingCables.pdf
|
|
16400
|
Thu Oct 14 09:28:46 2021 |
Yehonathan | Update | PSL | PMC unlocked |
PMC has been unlocked since ~ 2:30 AM. Seems like the PZT got saturated. I moved the DC output adjuster and the PMC locked immidiatly although with a low transmission of 0.62V (>0.7V is the usual case) and high REFL.
IMC locked immidiately but IFO seems to be completely misaligned. The beams on the AS monitor are moving quite alot syncronously. BS watchdog tripped. I enabled the coil outputs. Waiting for the RMS motion to relax...
Its not relaxing. RMS motion is still high. I disabled the coils again and reenabled them. This seems to have worked. Arms were locked quite easily but the ETMs oplevs were way off and the ASS couldn't get the TRX and TRY more than 0.7. I align the ETMs to center the oplev. I realign everything else and lock the arms. Maximium TR is still < 0.8.
|
16399
|
Wed Oct 13 15:36:38 2021 |
Hang | Update | Calibration | XARM OLTF |
We did a few quick XARM oltf measurements. We excited C1:LSC-ETMX_EXC with a broadband white noise upto 4 kHz. The timestamps for the measurements are: 1318199043 (start) - 1318199427 (end).
We will process the measurement to compute the cavity pole and analog filter poles & zeros later. |
Attachment 1: Screenshot_2021-10-13_15-32-16.png
|
|
16398
|
Wed Oct 13 11:25:14 2021 |
Anchal | Summary | CDS | Ran c1sus2 models in martian CDS. All good! |
Three extra steps (when adding new models, new FE):
- Chris pointed out that the sudo command in c1sus2 is giving error
sudo: unable to resolve host c1sus2
This error comes in when the computer could not figure out it's own hostname. Since FEs are network booted off the fb1, we need to update the /etc/hosts in /diskless/root everytime we add a new FE.
controls@fb1:~ 0$ sudo chroot /diskless/root
fb1:/ 0# sudo nano /etc/hosts
fb1:/ 0# exit
I added the following line in /etc/hosts file above:
192.168.113.92 c1sus2 c1sus2.martian
This resolved the issue of sudo giving error. Now, the rtcds make and install steps had no errors mentioned in their outputs.
- Another thing that needs to be done, as Koji pointed out, is to add the host and models in /etc/rtsystab in /diskless/root of fb:
controls@fb1:~ 0$ sudo chroot /diskless/root
fb1:/ 0# sudo nano /etc/rtsystab
fb1:/ 0# exit
I added the following lines in /etc/rtsystab file above:
c1sus2 c1x07 c1su2
This told rtcds what models would be available on c1sus2. Now rtcds list is displaying the right models:
controls@c1sus2:~ 0$ rtcds list
c1x07
c1su2
- The above steps are still not sufficient for the daqd_ processes to know about the new models. This part is supossed to happen automatically, but does not happen in our CDS apparently. So everytime there is a new model, we need to edit the file /opt/rtcds/caltech/c1/target/daqd/master and add following lines to it:
# Fast Data Channel lists
# c1sus2
/opt/rtcds/caltech/c1/chans/daq/C1X07.ini
/opt/rtcds/caltech/c1/chans/daq/C1SU2.ini
# test point lists
# c1sus2
/opt/rtcds/caltech/c1/target/gds/param/tpchn_c1x07.par
/opt/rtcds/caltech/c1/target/gds/param/tpchn_c1su2.par
I needed to restart the daqd_ processes in fb1 for them to notice these changes:
controls@fb1:~ 0$ sudo systemctl restart daqd_*
This finally lit up the status channels of DC in C1X07_GDS_TP.adl and C1SU2_GDS_TP.adl . However the channels C1:DAQ-DC0_C1X07_STATUS and C1:DAQ-DC0_C1SU2_STATUS both have values 0x2bad. This persists on restarting the models. I then just simply restarted teh mx_stream on c1sus2 and boom, it worked! (see attached all green screen, never seen before!)
So now Ian can work on testing the I/O chassis and we would be good to move c1sus2 FE and I/O chassis to 1Y3 after that. I've also done following extra changes:
- Updated CDS_FE_STATUS medm screen to show the new c1sus2 host.
- Updated global diag rest script to act on c1xo7 and c1su2 as well.
- Updated mxstream restart script to act on c1sus2 as well.
|
Attachment 1: CDS_screens_running.png
|
|
16397
|
Tue Oct 12 23:42:56 2021 |
Koji | Summary | CDS | Connected c1sus2 to martian network |
Don't you need to add the new hosts to /diskless/root/etc/rtsystab at fb1? --> There looks many elogs talking about editing "rtsystab".
controls@fb1:/diskless/root/etc 0$ cat rtsystab
#
# host list of control systems to run, starting with IOP
#
c1iscex c1x01 c1scx c1asx
c1sus c1x02 c1sus c1mcs c1rfm c1pem
c1ioo c1x03 c1ioo c1als c1omc
c1lsc c1x04 c1lsc c1ass c1oaf c1cal c1dnn c1daf
c1iscey c1x05 c1scy c1asy
#c1test c1x10 c1tst2
|
16396
|
Tue Oct 12 17:20:12 2021 |
Anchal | Summary | CDS | Connected c1sus2 to martian network |
I connected c1sus2 to the martian network by splitting the c1sim connection with a 5-way switch. I also ran another ethernet cable from the second port of c1sus2 to the DAQ network switch on 1X7.
Then I logged into chiara and added the following in chiara:/etc/dhcp/dhcpd.conf :
host c1sus2 {
hardware ethernet 00:25:90:06:69:C2;
fixed-address 192.168.113.92;
}
And following line in chiara:/var/lib/bind/martian.hosts :
c1sus2 A 192.168.113.92
Note that entires c1bhd is already added in these files, probably during some earlier testing by Gautam or Jon. Then I ran following to restart the dhcp server and nameserver:
~> sudo service bind9 reload
[sudo] password for controls:
* Reloading domain name service... bind9 [ OK ]
~> sudo service isc-dhcp-server restart
isc-dhcp-server stop/waiting
isc-dhcp-server start/running, process 25764
Now, As I switched on c1sus2 from front panel, it booted over network from fb1 like other FE machines and I was able to login to it by first logging to fb1 and then sshing to c1sus2.
Next, I copied the simulink models and the medm screens of c1x06, xc1x07, c1bhd, c1sus2 from the paths mentioned on this wiki page. I also copied the medm screens from chiara(clone):/opt/rtcds/caltech/c1/medm to martian network chiara in the appropriate places. I have placed the file /opt/rtcds/caltech/c1/medm/teststand_sitemap.adl which can be used to open sitemap for c1bhd and c1sus2 IOP and user models.
Then I logged into c1sus2 (via fb1) and did make, install, start procedure:
controls@c1sus2:~ 0$ rtcds make c1x07
buildd: /opt/rtcds/caltech/c1/rtbuild/release
### building c1x07...
Cleaning c1x07...
Done
Parsing the model c1x07...
Done
Building EPICS sequencers...
Done
Building front-end Linux kernel module c1x07...
Done
RCG source code directory:
/opt/rtcds/rtscore/branches/branch-3.4
The following files were used for this build:
/opt/rtcds/userapps/release/cds/c1/models/c1x07.mdl
Successfully compiled c1x07
***********************************************
Compile Warnings, found in c1x07_warnings.log:
***********************************************
***********************************************
controls@c1sus2:~ 0$ rtcds install c1x07
buildd: /opt/rtcds/caltech/c1/rtbuild/release
### installing c1x07...
Installing system=c1x07 site=caltech ifo=C1,c1
Installing /opt/rtcds/caltech/c1/chans/C1X07.txt
Installing /opt/rtcds/caltech/c1/target/c1x07/c1x07epics
Installing /opt/rtcds/caltech/c1/target/c1x07
Installing start and stop scripts
/opt/rtcds/caltech/c1/scripts/killc1x07
/opt/rtcds/caltech/c1/scripts/startc1x07
sudo: unable to resolve host c1sus2
Performing install-daq
Updating testpoint.par config file
/opt/rtcds/caltech/c1/target/gds/param/testpoint.par
/opt/rtcds/rtscore/branches/branch-3.4/src/epics/util/updateTestpointPar.pl -par_file=/opt/rtcds/caltech/c1/target/gds/param/archive/testpoint_211012_174226.par -gds_node=24 -site_letter=C -system=c1x07 -host=c1sus2
Installing GDS node 24 configuration file
/opt/rtcds/caltech/c1/target/gds/param/tpchn_c1x07.par
Installing auto-generated DAQ configuration file
/opt/rtcds/caltech/c1/chans/daq/C1X07.ini
Installing Epics MEDM screens
Running post-build script
safe.snap exists
controls@c1sus2:~ 0$ rtcds start c1x07
Cannot start/stop model 'c1x07' on host c1sus2.
controls@c1sus2:~ 4$ rtcds list
controls@c1sus2:~ 0$
One can see that even after making and installing, the model c1x07 is not listed as available models in rtcds list. Same is the case for c1sus2 as well. So I could not proceed with testing.
Good news is that nothing that I did affect the current CDS functioning. So we can probably do this testing safely from the main CDS setup. |
16395
|
Tue Oct 12 17:10:56 2021 |
Anchal | Summary | CDS | Some more information |
Chris pointed out some information displaying scripts, that show if the DAQ network is working or not. I thought it would be nice to log this information here as well.
controls@fb1:/opt/mx/bin 0$ ./mx_info
MX Version: 1.2.16
MX Build: controls@fb1:/opt/src/mx-1.2.16 Mon Aug 14 11:06:09 PDT 2017
1 Myrinet board installed.
The MX driver is configured to support a maximum of:
8 endpoints per NIC, 1024 NICs on the network, 32 NICs per host
===================================================================
Instance #0: 364.4 MHz LANai, PCI-E x8, 2 MB SRAM, on NUMA node 0
Status: Running, P0: Link Up
Network: Ethernet 10G
MAC Address: 00:60:dd:45:37:86
Product code: 10G-PCIE-8B-S
Part number: 09-04228
Serial number: 423340
Mapper: 00:60:dd:45:37:86, version = 0x00000000, configured
Mapped hosts: 3
ROUTE COUNT
INDEX MAC ADDRESS HOST NAME P0
----- ----------- --------- ---
0) 00:60:dd:45:37:86 fb1:0 1,0
1) 00:25:90:05:ab:47 c1bhd:0 1,0
2) 00:25:90:06:69:c3 c1sus2:0 1,0
controls@c1bhd:~ 1$ /opt/open-mx/bin/omx_info
Open-MX version 1.5.4
build: root@fb1:/opt/src/open-mx-1.5.4 Tue Aug 15 23:48:03 UTC 2017
Found 1 boards (32 max) supporting 32 endpoints each:
c1bhd:0 (board #0 name eth1 addr 00:25:90:05:ab:47)
managed by driver 'igb'
Peer table is ready, mapper is 00:60:dd:45:37:86
================================================
0) 00:25:90:05:ab:47 c1bhd:0
1) 00:60:dd:45:37:86 fb1:0
2) 00:25:90:06:69:c3 c1sus2:0
controls@c1sus2:~ 0$ /opt/open-mx/bin/omx_info
Open-MX version 1.5.4
build: root@fb1:/opt/src/open-mx-1.5.4 Tue Aug 15 23:48:03 UTC 2017
Found 1 boards (32 max) supporting 32 endpoints each:
c1sus2:0 (board #0 name eth1 addr 00:25:90:06:69:c3)
managed by driver 'igb'
Peer table is ready, mapper is 00:60:dd:45:37:86
================================================
0) 00:25:90:06:69:c3 c1sus2:0
1) 00:60:dd:45:37:86 fb1:0
2) 00:25:90:05:ab:47 c1bhd:0
These outputs prove that the framebuilder and the FEs are able to see each other in teh DAQ network.
Further, the error that we see when IOP model is started which crashes the mx_stream service on the FE machines (see 40m/16391) :
isendxxx failed with status Remote Endpoint Unreachable
This has been seen earlier when Jamie was troubleshooting the current fb1 in martian network in 40m/11655 in Oct, 2015. Unfortunately, I could not find what Jamie did over a year to fix this issue. |
16394
|
Tue Oct 12 16:39:52 2021 |
rana | Summary | SUS | PRM and BS Angular Actuation transfer function magnitude measurements |
should compare side by side with the ITM PRMI radar plots to see if there is a difference. How do your new plots compare with Gautam's plots of PRMI? |
16393
|
Tue Oct 12 11:32:54 2021 |
Yehonathan | Summary | SUS | PRM and BS Angular Actuation transfer function magnitude measurements |
Late submission (From Thursday 10/07):
I measured the PRMI sensing matrix to see if the BS and PRMI output matrices tweaking had any effect.
While doing so, I noticed I made a mistake in the analysis of the previous sensing matrix measurement. It seems that I have used the radar plot function with radians where degrees should have been used (the reason is that the azimuthal uncertainty looked crazy when I used degrees. I still don't know why this is the case with this measurement).
In any case, attachment 1 and 2 show the PRMI radar plots with the modified output matrices and and in the normal state, respectively.
It seems like the output matrix modification didn't do anything but REFL55 has good orthogonality. Problem gone?? |
Attachment 1: modified_output_matrices_radar_plots.png
|
|
Attachment 2: normal_output_matrices_radar_plots.png
|
|
16392
|
Mon Oct 11 18:29:35 2021 |
Anchal | Summary | CDS | Moving forward? |
The teststand has some non-trivial issue with Myrinet card (either software or hardware) which even teh experts are saying they don't remember how to fix it. CDS with mx was iin use more than a decade ago, so it is hard to find support for issues with it now and will be the same in future. We need to wrap up this test procedure one way or another now, so I have following two options moving forward:
Direct integration with main CDS and testing
- We can just connect the c1sus2 and c1bhd FE computers to martian network directly.
- We'll have to connect c1sus2 and c1bhd to the optical fiber subnetwork as well.
- On booting, they would get booted through the exisitng fb1 boot server which seems to work fine for the other 5 FE machines.
- We can update teh DHCP in chiara and reload it so that we can ssh into these FEs with host names.
- Hopefully, presence of these computers won't tank the existing CDS even if they themselves have any issues, as they have no shared memory with other models.
- If this works, we can do the loop back testing of I/O chassis using the main DAQ network and move on with our upgrade.
- If this does not work and causes any harm to exisitng CDS network, we can disconnect these computers and go back to existing CDS. Recently, our confidence on rebooting the CDS has increased with the robust performance as some legacy issues were fixed.
- We'll however, continue to use a CDS which is no more supported by the current LIGO CDS group.
Testing CDS upgrade on teststand
- From what I could gather, most of the hardware in I/O chassis that I could find, is still used in CDS of LLO and LHO, with their recent tests and documents using the same cards and PCBs.
- There might be some difference in the DAQ network setup that I need to confirm.
- I've summarised the current c1teststand hardware on this wiki page.
- If the latest CDS is backwards compatible with our hardware, we can test the new CDS in teh c1teststand setup without disrupting our main CDS. We'll have ample help and support for this upgrade from the current LIGO CDS group.
- We can do the loop back testing of the I/O chassis as well.
- If the upgrade is succesfull in the teststand without many hardware changes, we can upgrade the main CDS of 40m as well, as it has the same hardware as our teststand.
- Biggest plus point would be that out CDS will be up-to-date and we will be able to take help from CDS group if any trouble occurs.
So these are the two options we have. We should discuss which one to take in the mattermost chat or in upcoming meeting. |
16391
|
Mon Oct 11 17:31:25 2021 |
Anchal | Summary | CDS | Fixed mounting of mx devices in fb. daqd_dc is running now. |
I compared the fb1 in main network with the cloned fb1 and I found a crucial difference. The main fb1 where cds is running fine as mx devices mounted in /dev/ like mx0, mx1 upto mx7, mxctlm mxctlp, mxp0, mxp1 upto mxp7. The cloned fb does not have any of these mx devices mounted. I think this is where the issue was coming in from.
However, lspci | grep 'Myri' shows following output on both computers:
controls@fb1:/dev 0$ lspci | grep 'Myri'
02:00.0 Ethernet controller: MYRICOM Inc. Myri-10G Dual-Protocol NIC (rev 01)
Which means that the computer detects the card on PCie slot.
I tried to add this to /etc/rc.local to run this script at every boot, but it did not work. So for now, I'll just manually do this step everytime. Once the devices are loaded, we get:
controls@fb1:/etc 0$ ls /dev/*mx*
/dev/mx0 /dev/mx4 /dev/mxctl /dev/mxp2 /dev/mxp6 /dev/ptmx
/dev/mx1 /dev/mx5 /dev/mxctlp /dev/mxp3 /dev/mxp7
/dev/mx2 /dev/mx6 /dev/mxp0 /dev/mxp4 /dev/open-mx
/dev/mx3 /dev/mx7 /dev/mxp1 /dev/mxp5 /dev/open-mx-raw
The, restarting all daqd_ processes, I found that daqd_dc was running succesfully now. Here is the status:
controls@fb1:/etc 0$ sudo systemctl status daqd_* -l
● daqd_dc.service - Advanced LIGO RTS daqd data concentrator
Loaded: loaded (/etc/systemd/system/daqd_dc.service; enabled)
Active: active (running) since Mon 2021-10-11 17:48:00 PDT; 23min ago
Main PID: 2308 (daqd_dc_mx)
CGroup: /daqd.slice/daqd_dc.service
├─2308 /usr/bin/daqd_dc_mx -c /opt/rtcds/caltech/c1/target/daqd/daqdrc.dc
└─2370 caRepeater
Oct 11 17:48:07 fb1 daqd_dc_mx[2308]: mx receiver 006 thread priority error Operation not permitted[Mon Oct 11 17:48:06 2021]
Oct 11 17:48:07 fb1 daqd_dc_mx[2308]: mx receiver 005 thread put on CPU 0
Oct 11 17:48:07 fb1 daqd_dc_mx[2308]: [Mon Oct 11 17:48:06 2021] [Mon Oct 11 17:48:06 2021] mx receiver 006 thread put on CPU 0
Oct 11 17:48:07 fb1 daqd_dc_mx[2308]: mx receiver 007 thread put on CPU 0
Oct 11 17:48:07 fb1 daqd_dc_mx[2308]: [Mon Oct 11 17:48:06 2021] mx receiver 003 thread - label dqmx003 pid=2362
Oct 11 17:48:07 fb1 daqd_dc_mx[2308]: [Mon Oct 11 17:48:06 2021] mx receiver 003 thread priority error Operation not permitted
Oct 11 17:48:07 fb1 daqd_dc_mx[2308]: [Mon Oct 11 17:48:06 2021] mx receiver 003 thread put on CPU 0
Oct 11 17:48:07 fb1 daqd_dc_mx[2308]: warning:regcache incompatible with malloc
Oct 11 17:48:07 fb1 daqd_dc_mx[2308]: [Mon Oct 11 17:48:06 2021] EDCU has 410 channels configured; first=0
Oct 11 17:49:06 fb1 daqd_dc_mx[2308]: [Mon Oct 11 17:49:06 2021] ->4: clear crc
● daqd_fw.service - Advanced LIGO RTS daqd frame writer
Loaded: loaded (/etc/systemd/system/daqd_fw.service; enabled)
Active: active (running) since Mon 2021-10-11 17:48:01 PDT; 23min ago
Main PID: 2318 (daqd_fw)
CGroup: /daqd.slice/daqd_fw.service
└─2318 /usr/bin/daqd_fw -c /opt/rtcds/caltech/c1/target/daqd/daqdrc.fw
Oct 11 17:48:09 fb1 daqd_fw[2318]: [Mon Oct 11 17:48:09 2021] [Mon Oct 11 17:48:09 2021] Producer thread - label dqproddbg pid=2440
Oct 11 17:48:09 fb1 daqd_fw[2318]: Producer crc thread priority error Operation not permitted
Oct 11 17:48:09 fb1 daqd_fw[2318]: [Mon Oct 11 17:48:09 2021] [Mon Oct 11 17:48:09 2021] Producer crc thread put on CPU 0
Oct 11 17:48:09 fb1 daqd_fw[2318]: Producer thread priority error Operation not permitted
Oct 11 17:48:09 fb1 daqd_fw[2318]: [Mon Oct 11 17:48:09 2021] Producer thread put on CPU 0
Oct 11 17:48:09 fb1 daqd_fw[2318]: [Mon Oct 11 17:48:09 2021] Producer thread - label dqprod pid=2434
Oct 11 17:48:09 fb1 daqd_fw[2318]: [Mon Oct 11 17:48:09 2021] Producer thread priority error Operation not permitted
Oct 11 17:48:09 fb1 daqd_fw[2318]: [Mon Oct 11 17:48:09 2021] Producer thread put on CPU 0
Oct 11 17:48:10 fb1 daqd_fw[2318]: [Mon Oct 11 17:48:10 2021] Minute trender made GPS time correction; gps=1318034906; gps%60=26
Oct 11 17:49:09 fb1 daqd_fw[2318]: [Mon Oct 11 17:49:09 2021] ->3: clear crc
● daqd_rcv.service - Advanced LIGO RTS daqd testpoint receiver
Loaded: loaded (/etc/systemd/system/daqd_rcv.service; enabled)
Active: active (running) since Mon 2021-10-11 17:48:00 PDT; 23min ago
Main PID: 2311 (daqd_rcv)
CGroup: /daqd.slice/daqd_rcv.service
└─2311 /usr/bin/daqd_rcv -c /opt/rtcds/caltech/c1/target/daqd/daqdrc.rcv
Oct 11 17:50:21 fb1 daqd_rcv[2311]: Creating C1:DAQ-NDS0_C1X07_CRC_SUM
Oct 11 17:50:21 fb1 daqd_rcv[2311]: Creating C1:DAQ-NDS0_C1BHD_STATUS
Oct 11 17:50:21 fb1 daqd_rcv[2311]: Creating C1:DAQ-NDS0_C1BHD_CRC_CPS
Oct 11 17:50:21 fb1 daqd_rcv[2311]: Creating C1:DAQ-NDS0_C1BHD_CRC_SUM
Oct 11 17:50:21 fb1 daqd_rcv[2311]: Creating C1:DAQ-NDS0_C1SU2_STATUS
Oct 11 17:50:21 fb1 daqd_rcv[2311]: Creating C1:DAQ-NDS0_C1SU2_CRC_CPS
Oct 11 17:50:21 fb1 daqd_rcv[2311]: Creating C1:DAQ-NDS0_C1SU2_CRC_SUM
Oct 11 17:50:21 fb1 daqd_rcv[2311]: Creating C1:DAQ-NDS0_C1OM[Mon Oct 11 17:50:21 2021] Epics server started
Oct 11 17:50:24 fb1 daqd_rcv[2311]: [Mon Oct 11 17:50:24 2021] Minute trender made GPS time correction; gps=1318035040; gps%120=40
Oct 11 17:51:21 fb1 daqd_rcv[2311]: [Mon Oct 11 17:51:21 2021] ->3: clear crc
Now, even before starting teh FE models, I see DC status as ox2bad in the CDS screens of the IOP and user models. The mx_stream service remains in a failed state at teh FE machines and remain the same even after restarting the service.
controls@c1sus2:~ 0$ sudo systemctl status mx_stream -l
● mx_stream.service - Advanced LIGO RTS front end mx stream
Loaded: loaded (/etc/systemd/system/mx_stream.service; enabled)
Active: failed (Result: exit-code) since Mon 2021-10-11 17:50:26 PDT; 15min ago
Process: 382 ExecStart=/etc/mx_stream_exec (code=exited, status=1/FAILURE)
Main PID: 382 (code=exited, status=1/FAILURE)
Oct 11 17:50:25 c1sus2 systemd[1]: Starting Advanced LIGO RTS front end mx stream...
Oct 11 17:50:25 c1sus2 systemd[1]: Started Advanced LIGO RTS front end mx stream.
Oct 11 17:50:25 c1sus2 mx_stream_exec[382]: Failed to open endpoint Not initialized
Oct 11 17:50:26 c1sus2 systemd[1]: mx_stream.service: main process exited, code=exited, status=1/FAILURE
Oct 11 17:50:26 c1sus2 systemd[1]: Unit mx_stream.service entered failed state.
But if I restart the mx_stream service before starting the rtcds models, the mx-stream service starts succesfully:
controls@c1sus2:~ 0$ sudo systemctl restart mx_stream
controls@c1sus2:~ 0$ sudo systemctl status mx_stream -l
● mx_stream.service - Advanced LIGO RTS front end mx stream
Loaded: loaded (/etc/systemd/system/mx_stream.service; enabled)
Active: active (running) since Mon 2021-10-11 18:14:13 PDT; 25s ago
Main PID: 1337 (mx_stream)
CGroup: /system.slice/mx_stream.service
└─1337 /usr/bin/mx_stream -e 0 -r 0 -w 0 -W 0 -s c1x07 c1su2 -d fb1:0
Oct 11 18:14:13 c1sus2 systemd[1]: Starting Advanced LIGO RTS front end mx stream...
Oct 11 18:14:13 c1sus2 systemd[1]: Started Advanced LIGO RTS front end mx stream.
Oct 11 18:14:13 c1sus2 mx_stream_exec[1337]: send len = 263596
Oct 11 18:14:13 c1sus2 mx_stream_exec[1337]: Connection Made
However, the DC status on CDS screens still show 0x2bad. As soon as I start the rtcds model c1x07 (the IOP model for c1sus2), the mx_stream service fails:
controls@c1sus2:~ 0$ sudo systemctl status mx_stream -l
● mx_stream.service - Advanced LIGO RTS front end mx stream
Loaded: loaded (/etc/systemd/system/mx_stream.service; enabled)
Active: failed (Result: exit-code) since Mon 2021-10-11 18:18:03 PDT; 27s ago
Process: 1337 ExecStart=/etc/mx_stream_exec (code=exited, status=1/FAILURE)
Main PID: 1337 (code=exited, status=1/FAILURE)
Oct 11 18:14:13 c1sus2 systemd[1]: Starting Advanced LIGO RTS front end mx stream...
Oct 11 18:14:13 c1sus2 systemd[1]: Started Advanced LIGO RTS front end mx stream.
Oct 11 18:14:13 c1sus2 mx_stream_exec[1337]: send len = 263596
Oct 11 18:14:13 c1sus2 mx_stream_exec[1337]: Connection Made
Oct 11 18:18:03 c1sus2 mx_stream_exec[1337]: isendxxx failed with status Remote Endpoint Unreachable
Oct 11 18:18:03 c1sus2 mx_stream_exec[1337]: disconnected from the sender
Oct 11 18:18:03 c1sus2 mx_stream_exec[1337]: c1x07_daq mmapped address is 0x7fe3620c3000
Oct 11 18:18:03 c1sus2 mx_stream_exec[1337]: c1su2_daq mmapped address is 0x7fe35e0c3000
Oct 11 18:18:03 c1sus2 systemd[1]: mx_stream.service: main process exited, code=exited, status=1/FAILURE
Oct 11 18:18:03 c1sus2 systemd[1]: Unit mx_stream.service entered failed state.
This shows that the start of rtcds model, causes the fail in mx_stream, possibly due to inability of finding the endpoint on fb1. I've again reached to the edge of my knowledge here. Maybe the fiber optic connection between fb and the network switch that connects to FE is bad, or the connection between switch and FEs is bad.
But we are just one step away from making this work.
|
16390
|
Mon Oct 11 13:59:47 2021 |
Hang | Update | SUS | More PRM L2P measurements |
We report here the analysis results for the measurements done in elog:16388.
Figs. 1 & 2 are respectively measurements of the white noise excitation and the optimized excitation. The shaded region corresponds to the 1-sigma uncertainty at each frequency bin. By eyes, one can already see that the constraints on the phase in the 0.6-1 Hz band are much tighter in the optimized case than in the white noise case.
We found the transfer function was best described by two real poles + one pair of complex poles (i.e., resonance) + one pair of complex zeros in the right-half plane (non-minimum phase delay). The measurement in fact suggested a right-hand pole somewhere between 0.05-0.1 Hz which cannot be right. For now, I just manually flipped the sign of this lowest frequency pole to the left-hand side. However, this introduced some systematic deviation in the phase in the 0.3-0.5 Hz band where our coherence was still good. Therefore, a caveat is that our model with 7 free parameters (4 poles + 2 zeros + 1 gain as one would expect for an ideal signal-stage L2P TF) might not sufficiently capture the entire physics.
In Fig. 3 we showed the comparison of the two sets of measurements together with the predictions based on the Fisher matrix. Here the color gray is for the white-noise excitation and olive is for the optimized excitation. The solid and dotted contours are respectively the 1-sigma and 3-sigma regions from the Fisher calculation, and crosses are maximum likelihood estimations of each measurement (though the scipy optimizer might not find the true maximum).
Note that the mean values don't match in the two sets of measurements, suggesting potential bias or other systematics exists in the current measurement. Moreover, there could be multiple local maxima in the likelihood in this high-D parameter space (not surprising). For example, one could reduce the resonant Q but enhance the overall gain to keep the shoulder of a resonance having the same amplitude. However, this correlation is not explicit in the Fisher matrix (first-order derivatives of the TF, i.e., local gradients) as it does not show up in the error ellipse.
In Fig. 4 we show the further optimized excitation for the next round of measurements. Here the cyan and olive traces are obtained assuming different values of the "true" physical parameter, yet the overall shapes of the two are quite similar, and are close to the optimized excitation spectrum we already used in elog:16388.
|
Attachment 1: prm_l2p_tf_meas_white.pdf
|
|
Attachment 2: prm_l2p_tf_meas_opt.pdf
|
|
Attachment 3: prm_l2p_fisher_vs_data_white_vs_opt.pdf
|
|
Attachment 4: prm_l2p_Pxx_evol_v2.pdf
|
|
16389
|
Mon Oct 11 11:13:04 2021 |
rana | Update | SUS | More PRM L2P measurements |
For the oplev, there are DQ channels you can use so that its possible to look back in the past for long measurements. They have names like PERROR |
16388
|
Fri Oct 8 17:33:13 2021 |
Hang | Update | SUS | More PRM L2P measurements |
[Raj, Hang]
We did some more measurements on the PRM L2P TF.
We tried to compare the parameter estimation uncertainties of white vs. optimal excitation. We drove C1:SUS-PRM_LSC_EXC with "Normal" excitation and digital gain of 700.
For the white noise exciation, we simply put a butter("LowPass",4,10) filter to select out the <10 Hz band.
For the optimal exciation, we use butter("BandPass",6,0.3,1.6) gain(3) notch(1,20,8) to approximate the spectral shape reported in elog:16384. We tried to use awg.ArbitraryLoop yet this function seems to have some bugs and didn't run correctly; an issue has been submitted to the gitlab repo with more details. We also noticed that in elog:16384, the pitch motion should be read out from C1:SUS-PRM_OL_PIT_IN1 instead of the OUT channel, as there are some extra filters between IN1 and OUT. Consequently, the exact optimal exciation should be revisited, yet we think the main result should not be altered significantly.
While a more detail analysis will be done later offline, we post in the attached plot a comparison between the white (blue) vs optimal (red) excitation. Note in this case, we kept the total force applied to the PRM the same (as the RMS level matches).
Under this simple case, the optimal excitation appears reasonable in two folds.
First, the optimization tries to concentrate the power around the resonance. We would naturally expect that near the resonance, we would get more Fisher information, as the phase changes the fastest there (i.e., large derivatives in the TF).
Second, while we move the power in the >2 Hz band to the 0.3-2 Hz band, from the coherence plot we see that we don't lose any information in the > 2 Hz region. Indeed, even with the original white excitation, the coherence is low and the > 2 Hz region would not be informative. Therefore, it seems reasonable to give up this band so that we can gain more information from locations where we have meaningful coherence. |
Attachment 1: Screenshot_2021-10-08_17-30-52.png
|
|
16387
|
Thu Oct 7 02:04:19 2021 |
Koji | Update | Electronics | Satellite amp adapter chassis |
The 4 units of Satellite Amp Adapter were done:
- The ears were fixed with the screws
- The handles were attached (The stock of the handles is low)
- The boards are now supported by plastic stand-offs. (The chassis were drilled)
- The front and rear panels were fixed to the chassis
- The front and rear connectors were fixed with the low profile 4-40 stand-off screws (3M 3341-1S)
|
Attachment 1: P_20211006_205044.jpg
|
|
16386
|
Wed Oct 6 16:31:02 2021 |
Tega | Update | Electronics | Sat Amp modifications |
[Tega, Koji]
(S2100737) - Debugging showed that the opamp, AD822ARZ, for PD2 circuit was not working as expected so we replaced with a spare and this fixed the problem. Somehow, the PD1 circuit no longer presents any issues, so everything is now fine with the unit.
(S2100741) - All good.
Quote: |
Trying to finish 2 more Sat Amp units so that we have the 7 units needed for the X-arm install.
S2100736 - All good
S2100737 - This unit presented with an issue on the PD1 circuit of channel 1-4 PCB where the voltage reading on TP6, TP7 and TP8 are -15.1V, -14.2V, and +14.7V respectively, instead of ~0V. The unit also has an issue on the PD2 circuit of channel 1-4 PCB because the voltage reading on TP7 and TP8 are -14.2V, and +14.25V respectively, instead of ~0V.
|
|
16385
|
Wed Oct 6 15:39:29 2021 |
Anchal | Summary | SUS | PRM and BS Angular Actuation transfer function magnitude measurements |
Note that your tests were done with the output matrix for BS and PRM in the compensated state as done in 40m/16374. The changes made there were supposed to clear out any coil actuation imbalance in the angular degrees of freedom. |
16384
|
Wed Oct 6 15:04:36 2021 |
Hang | Update | SUS | PRM L2P TF measurement & Fisher matrix analysis |
[Paco, Hang]
Yesterday afternoon Paco and I measured the PRM L2P transfer function. We drove C1:SUS-PRM_LSC_EXC with a white noise in the 0-10 Hz band (effectively a white, longitudinal force applied to the suspension) and read out the pitch response in C1:SUS-PRM_OL_PIT_OUT. The local damping was left on during the measurement. Each FFT segment in our measurement is 32 sec and we used 8 non-overlapping segments for each measurement. The empirically determined results are also compared with the Fisher matrix estimation (similar to elog:16373).
Results:
Fig. 1 shows one example of the measured L2P transfer function. The gray traces are measurement data and shaded region the corresponding uncertainty. The olive trace is the best fit model.
Note that for a single-stage suspension, the ideal L2P TF should have two zeros at DC and two pairs of complex poles for the length and pitch resonances, respectively. We found the two resonances at around 1 Hz from the fitting as expected. However, the zeros were not at DC as the ideal, theoretical model suggested. Instead, we found a pair of right-half plane zeros in order to explain the measurement results. If we cast such a pair of right-half plane zeros into (f, Q) pair, it means a negative value of Q. This means the system does not have the minimum phase delay and suggests some dirty cross-coupling exists, which might not be surprising.
Fig. 2 compares the distribution of the fitting results for 4 different measurements (4 red crosses) and the analytical error estimation obtained using the Fisher matrix (the gray contours; the inner one is the 1-sigma region and the outer one the 3-sigma region). The Fisher matrix appears to underestimate the scattering from this experiment, yet it does capture the correlation between different parameters (the frequencies and quality factors of the two resonances).
One caveat though is that the fitting routine is not especially robust. We used the vectfit routine w/ human intervening to get some initial guesses of the model. We then used a standard scipy least-sq routine to find the maximal likelihood estimator of the restricted model (with fixed number of zeros and poles; here 2 complex zeros and 4 complex poles). The initial guess for the scipy routine was obtained from the vectfit model.
Fig. 3 shows how we may shape our excitation PSD to maximize the Fisher information while keeping the RMS force applied to the PRM suspension fixed. In this case the result is very intuitive. We simply concentrate our drive around the resonance at ~ 1 Hz, focusing on locations where we initially have good SNR. So at least code is not suggesting something crazy...
Fig. 4 then shows how the new uncertainty (3-sigma contours) should change as we optimize our excitation. Basically one iteration (from gray to olive) is sufficient here.
We will find a time very recently to repeat the measurement with the optimized injection spectrum. |
Attachment 1: prm_l2p_tf_meas.pdf
|
|
Attachment 2: prm_l2p_fisher_vs_data.pdf
|
|
Attachment 3: prm_l2p_Pxx_evol.pdf
|
|
Attachment 4: prm_l2p_fisher_evol.pdf
|
|
16383
|
Tue Oct 5 20:04:22 2021 |
Paco | Summary | SUS | PRM and BS Angular Actuation transfer function magnitude measurements |
[Paco, Rana]
We had a look at the BS actuation. Along the way we created a couple of issues that we fixed. A summary is below.
- First, we locked MICH. While doing this, we used the /users/Templates/ndscope/LSC/MICH.yml ndscope template to monitor some channels. I edited the yaml file to look at C1:LSC-ASDC_OUT_DQ instead of the REFL_DC. Rana pointed out that the C1:LSC-MICH_OUT_DQ (MICH control point) had a big range (~ 5000 counts rms) and this should not be like that.
- We tried to investigate the aforementioned thing by looking at the whitening / uwhitening filters but all the slow epics channels where "white" on the medm screen. Looking under CDS/slow channel monitors, we realized that both c1iscaux and c1auxey were weird, so we tried telnet to c1iscaux without success. Therefore, we followed the recommended wiki procedure of hard rebooting this machine. While inside the lab and looking for this machine, we touched things around the 'rfpd' rack and once we were back in the control room, we couldn't see any light on the AS port camera. But the whitening filter medm screens were back up.
- While rana ssh'd into c1auxey to investigate about its status, and burtrestored the c1iscaux channels, we looked at trends to figure out if anything had changed (for example TT1 or TT2) but this wasn't the case. We decided to go back inside to check the actual REFL beams and noticed it was grossly misaligned (clipping)... so we blamed it on the TTs and again, went around and moved some stuff around the 'rfpd' rack. We didn't really connect or disconnect anything, but once we were back in the control room, light was coming from the AS port again. This is a weird mystery and we should systematically try to repeat this and fix the actual issue.
- We restored the MICH, and returned to BS actuation problems. Here, we essentially devised a scheme to inject noise at 310.97 Hz and 313.74. The choice is twofold, first it lies outside the MICH loop UGF (~150 Hz), and second, it matches the sensing matrix OSC frequencies, so it's more appropriate for a comparison.
- We injected two lines using the BS SUS LOCKIN1 and LOCKIN2 oscilators so we can probe two coils at once, with the LSC loop closed, and read back using the C1:LSC-MICH_IN1_DQ channel. We excited with an amplitude of 1234.0 counts and 1254 counts respectively (to match the ~ 2 % difference in frequency) and noted that the magnitude response in UR was 10% larger than UL, LL, and LR which were close to each other at the 2% level.
[Paco]
After rana left, I did a second pass at the BS actuation. I took TF measurements at the oscilator frequencies noted above using diaggui, and summarize the results below:
TF |
UL (310.97 Hz) |
UR (313.74 Hz) |
LL (310.97 Hz) |
LR (313.74 Hz) |
Magnitude (dB) |
93.20 |
92.20 |
94.27 |
93.85 |
Phase (deg) |
-128.3 |
-127.9 |
-128.4 |
-127.5 |
This procedure should be done with PRM as well and using the PRCL instead of MICH. |
16382
|
Tue Oct 5 18:00:53 2021 |
Anchal | Summary | CDS | c1teststand time synchronization working now |
Today I got a new router that I used to connect the c1teststand, fb1 and chiara. I was able to see internet access in c1teststand and fb1, but not in chiara. I'm not sure why that is the case.
The good news is that the ntp server on fb1(clone) is working fine now and both FE computers, c1bhd and c1sus2 are succesfully synchronized to the fb1(clone) ntpserver. This resolves any possible timing issues in this DAQ network.
On running the IOP and user models however, I see the same errors are mentioned in 40m/16372. Something to do with:
Oct 06 00:47:56 c1sus2 mx_stream_exec[21796]: OMX: Failed to find peer index of board 00:00:00:00:00:00 (Peer Not Found in the Table)
Oct 06 00:47:56 c1sus2 mx_stream_exec[21796]: mx_connect failed Nic ID not Found in Peer Table
Oct 06 00:47:56 c1sus2 mx_stream_exec[21796]: c1x07_daq mmapped address is 0x7fa4819cc000
Oct 06 00:47:56 c1sus2 mx_stream_exec[21796]: c1su2_daq mmapped address is 0x7fa47d9cc000
Thu Oct 7 17:04:31 2021
I fixed the issue of chiara not getting internet. Now c1teststand, fb1 and chiara, all have internet connections. It was the issue of default gateway and interface and findiing the DNS. I have found the correct settings now. |
16381
|
Tue Oct 5 17:58:52 2021 |
Anchal | Summary | CDS | c1teststand problems summary |
open-mx service is running successfully on the fb1(clone), c1bhd and c1sus.
Quote: |
I don't know anything about mx/open-mx, but you also need open-mx,don't you?
|
|
16380
|
Tue Oct 5 17:01:20 2021 |
Koji | Update | Electronics | Sat Amp modifications |
Make sure the inputs for the PD amps are open. This is the current amplifier and we want to leave the input pins open for the test of this circuit.
TP6 is the first stage of the amps (TIA). So this stage has the issue. Usual check if the power is properly supplied / if the pins are properly connected/isolated / If the opamp is alive or not.
For TP8, if TP8 get railed. TP5 and TP7 are going to be railed too. Is that the case, if so, check this whitening stage in the same way as above.
If the problem is only in the TP5 and/or TP7 it is the differential driver issue. Check the final stage as above. Replacing the opamp could help.
|
16379
|
Mon Oct 4 21:58:17 2021 |
Tega | Update | Electronics | Sat Amp modifications |
Trying to finish 2 more Sat Amp units so that we have the 7 units needed for the X-arm install.
S2100736 - All good
S2100737 - This unit presented with an issue on the PD1 circuit of channel 1-4 PCB where the voltage reading on TP6, TP7 and TP8 are -15.1V, -14.2V, and +14.7V respectively, instead of ~0V. The unit also has an issue on the PD2 circuit of channel 1-4 PCB because the voltage reading on TP7 and TP8 are -14.2V, and +14.25V respectively, instead of ~0V.
|
16378
|
Mon Oct 4 20:46:08 2021 |
Koji | Update | Electronics | Satellite amp box adapters |
Thanks. You should be able to find the chassis-related hardware on the left side of the benchtop drawers at the middle workbench.
Hardware: The special low profile 4-40 standoff screw / 1U handles / screws and washers for the chassis / flat-top screws for chassis panels and lids |
16377
|
Mon Oct 4 18:35:12 2021 |
Paco | Update | Electronics | Satellite amp box adapters |
[Paco]
I have finished assembling the 1U adapters from 8 to 5 DB9 conn. for the satellite amp boxes. One thing I had to "hack" was the corners of the front panel end of the PCB. Because the PCB was a bit too wide, it wasn't really flush against the front panel (see Attachment #1), so I just filed the corners by ~ 3 mm and covered with kapton tape to prevent contact between ground planes and the chassis. After this, I made DB9 cables, connected everything in place and attached to the rear panel (Attachment #2). Four units are resting near the CAD machine (next to the bench area), see Attachment #3. |
Attachment 1: pcb_no_flush.jpg
|
|
Attachment 2: 1U_assembly.jpg
|
|
Attachment 3: fourunits.jpg
|
|
16376
|
Mon Oct 4 18:00:16 2021 |
Koji | Summary | CDS | c1teststand problems summary |
I don't know anything about mx/open-mx, but you also need open-mx,don't you?
controls@c1ioo:~ 0$ systemctl status *mx*
● open-mx.service - LSB: starts Open-MX driver
Loaded: loaded (/etc/init.d/open-mx)
Active: active (running) since Wed 2021-09-22 11:54:39 PDT; 1 weeks 5 days ago
Process: 470 ExecStart=/etc/init.d/open-mx start (code=exited, status=0/SUCCESS)
CGroup: /system.slice/open-mx.service
└─620 /opt/3.2.88-csp/open-mx-1.5.4/bin/fma -d
● mx_stream.service - Advanced LIGO RTS front end mx stream
Loaded: loaded (/etc/systemd/system/mx_stream.service; enabled)
Active: active (running) since Wed 2021-09-22 12:08:00 PDT; 1 weeks 5 days ago
Main PID: 5785 (mx_stream)
CGroup: /system.slice/mx_stream.service
└─5785 /usr/bin/mx_stream -e 0 -r 0 -w 0 -W 0 -s c1x03 c1ioo c1als c1omc -d fb1:0
|
16375
|
Mon Oct 4 16:10:09 2021 |
rana | Summary | SUS | PRM and BS Angular Actuation transfer function magnitude measurements |
not sure that this is necessary. If you look at teh previous entries Gautam made on this topic, it is clear that the BS/PRM PRMI matrix is snafu, whereas the ITM PRMI matrix is not.
Is it possible that the ~5% coil imbalance of the BS/PRM can explain the observed sensing matrix? If not, then there is no need to balance these coils. |
16374
|
Mon Oct 4 16:00:57 2021 |
Yehonathan | Summary | SUS | PRM and BS Angular Actuation transfer function magnitude measurements |
{Yehonathan, Anchel}
In an attempt to fix the actuation of the PRMI DOFs we set to modify the output matrix of the BS and PRM such that the response of the coils will be similar to each other as much as possible.
To do so, we used the responses at a single frequency from the previous measurement to infer the output matrix coefficients that will equilize the OpLev responses (arbitrarily making the LL coil as a reference). This corrected the imbalance in BS almost completely while it didn't really work for PRM (see attachment 1).
The new output matrices are shown in attachment 2-3. |
Attachment 1: BS_PRM_ANG_ACT_TF_20211004.pdf
|
|
Attachment 2: BS_out_mat_20211004.txt
|
9.839999999999999858e-01 8.965770586285104482e-01 9.486710352885977526e-01 3.099999999999999978e-01
1.016000000000000014e+00 9.750242104232501594e-01 -9.291967546765563801e-01 3.099999999999999978e-01
9.839999999999999858e-01 -1.086765190351774768e+00 1.009798093279114628e+00 3.099999999999999978e-01
1.016000000000000014e+00 -1.031706735496689786e+00 -1.103142995587099939e+00 3.099999999999999978e-01
0.000000000000000000e+00 0.000000000000000000e+00 0.000000000000000000e+00 1.000000000000000000e+00
|
Attachment 3: PRM_out_mat_20211004.txt
|
1.000000000000000000e+00 1.033455230230304611e+00 9.844796282226820905e-01 0.000000000000000000e+00
1.000000000000000000e+00 9.342329554807877745e-01 -1.021296201828568506e+00 0.000000000000000000e+00
1.000000000000000000e+00 -1.009214777246558503e+00 9.965113815550634691e-01 0.000000000000000000e+00
1.000000000000000000e+00 -1.020129700278567197e+00 -9.973560027273553619e-01 0.000000000000000000e+00
0.000000000000000000e+00 0.000000000000000000e+00 0.000000000000000000e+00 1.000000000000000000e+00
|
16373
|
Mon Oct 4 15:50:31 2021 |
Hang | Update | Calibration | Fisher matrix estimation on XARM parameters |
[Anchal, Hang]
What: Anchal and I measured the XARM OLTF last Thursday.
Goal: 1. measure the 2 zeros and 2 poles in the analog whitening filter, and potentially constrain the cavity pole and an overall gain.
2. Compare the parameter distribution obtained from measurements and that estimated analytically from the Fisher matrix calculation.
3. Obtain the optimized excitation spectrum for future measurements.
How: we inject at C1:SUS-ETMX_LSC_EXC so that each digital count should be directly proportional to the force applied to the suspension. We read out the signal at C1:SUS-ETMX_LSC_OUT_DQ. We use an approximately white excitation in the 50-300 Hz band, and intentionally choose the coherence to be only slightly above 0.9 so that we can get some statistical error to be compared with the Fisher matrix's prediction. For each measurement, we use a bandwidth of 0.25 Hz and 10 averages (no overlapping between adjacent segments).
The 2 zeros and 2 poles in the analog whitening filter and an overall gain are treated as free parameters to be fitted, while the rest are taken from the model by Anchal and Paco (elog:16363). The optical response of the arm cavity seems missing in that model, and thus we additionally include a real pole (for the cavity pole) in the model we fit. Thus in total, our model has 6 free parameters, 2 zeros, 3 poles, and 1 overall gain.
The analysis codes are pushed to the 40m/sysID repo.
===========================================================
Results:
Fig. 1 shows one measurement. The gray trace is the data and the olive one is the maximum likelihood estimation. The uncertainty for each frequency bin is shown in the shaded region. Note that the SNR is related to the coherence as
SNR^2 = [coherence / (1-coherence)] * (# of average),
and for a complex TF written as G = A * exp[1j*Phi], one can show the uncertainty is given by
\Delta A / A = 1/SNR, \Delta \Phi = 1/SNR [rad].
Fig. 2. The gray contours show the 1- and 2-sigma levels of the model parameters using the Fisher matrix calculation. We repeated the measurement shown in Fig. 1 three times, and the best-fit parameters for each measurement are indicated in the red-crosses. Although we only did a small number of experiments, the amount of scattering is consistent with the Fisher matrix's prediction, giving us some confidence in our analytical calculation.
One thing to note though is that in order to fit the measured data, we would need an additional pole at around 1,500 Hz. This seems a bit low for the cavity pole frequency. For aLIGO w/ 4km arms, the single-arm pole is about 40-50 Hz. The arm is 100 times shorter here and I would naively expect the cavity pole to be at 3k-4k Hz if the test masses are similar.
Fig. 3. We then follow the algorithm outlined in Pintelon & Schoukens, sec. 5.4.2.2, to calculate how we should change the excitation spectrum. Note that here we are fixing the rms of the force applied to the suspension constant.
Fig. 4 then shows how the expected error changes as we optimize the excitation. It seems in this case a white-ish excitation is already decent (as the TF itself is quite flat in the range of interest), and we only get some mild improvement as we iterate the excitation spectra (note we use the color gray, olive, and purple for the results after the 0th, 1st, and 2nd iteration; same color-coding as in Fig. 3).
|
Attachment 1: tf_meas.pdf
|
|
Attachment 2: fisher_est_vs_data.pdf
|
|
Attachment 3: Pxx_evol.pdf
|
|
Attachment 4: fisher_evol.pdf
|
|
16372
|
Mon Oct 4 11:05:44 2021 |
Anchal | Summary | CDS | c1teststand problems summary |
[Anchal, Paco]
We tried to fix the ntp synchronization in c1teststand today by repeating the steps listed in 40m/16302. Even though teh cloned fb1 now has the exact same package version, conf & service files, and status, the FE machines (c1bhd and c1sus2) fail to sync to the time. the timedatectl shows the same stauts 'Idle'. We also, dug bit deeper into the error messages of daq_dc on cloned fb1 and mx_stream on FE machines and have some error messages to report here.
Attempt on fixing the ntp
- We copied the ntp package version 1:4.2.6 deb file from /var/cache/apt/archives/ntp_1%3a4.2.6.p5+dfsg-7+deb8u3_amd64.deb on the martian fb1 to the cloned fb1 and ran.
controls@fb1:~ 0$ sudo dbpg -i ntp_1%3a4.2.6.p5+dfsg-7+deb8u3_amd64.deb
- We got error messages about missing dependencies of libopts25 and libssl1.1. We downloaded oldoldstable jessie versions of these packages from here and here. We ensured that these versions are higher than the required versions for ntp. We installed them with:
controls@fb1:~ 0$ sudo dbpg -i libopts25_5.18.12-3_amd64.deb
controls@fb1:~ 0$ sudo dbpg -i libssl1.1_1.1.0l-1~deb9u4_amd64.deb
- Then we installed the ntp package as described above. It asked us if we want to keep the configuration file, we pressed Y.
- However, we decided to make the configuration and service files exactly same as martian fb1 to make it same in cloned fb1. We copied /etc/ntp.conf and /etc/systemd/system/ntp.service files from martian fb1 to cloned fb1 in the same positions. Then we enabled ntp, reloaded the daemon, and restarted ntp service:
controls@fb1:~ 0$ sudo systemctl enable ntp
controls@fb1:~ 0$ sudo systemctl daemon-reload
controls@fb1:~ 0$ sudo systemctl restart ntp
- But ofcourse, since fb1 doesn't have internet access, we got some errors in status of the ntp.service:
controls@fb1:~ 0$ sudo systemctl status ntp
● ntp.service - NTP daemon (custom service)
Loaded: loaded (/etc/systemd/system/ntp.service; enabled)
Active: active (running) since Mon 2021-10-04 17:12:58 UTC; 1h 15min ago
Main PID: 26807 (code=exited, status=0/SUCCESS)
CGroup: /system.slice/ntp.service
├─30408 /usr/sbin/ntpd -p /var/run/ntpd.pid -g -u 105:107
└─30525 /usr/sbin/ntpd -p /var/run/ntpd.pid -g -u 105:107
Oct 04 17:48:42 fb1 ntpd_intres[30525]: host name not found: 2.debian.pool.ntp.org
Oct 04 17:48:52 fb1 ntpd_intres[30525]: host name not found: 3.debian.pool.ntp.org
Oct 04 18:05:05 fb1 ntpd_intres[30525]: host name not found: 0.debian.pool.ntp.org
Oct 04 18:05:15 fb1 ntpd_intres[30525]: host name not found: 1.debian.pool.ntp.org
Oct 04 18:05:25 fb1 ntpd_intres[30525]: host name not found: 2.debian.pool.ntp.org
Oct 04 18:05:35 fb1 ntpd_intres[30525]: host name not found: 3.debian.pool.ntp.org
Oct 04 18:21:48 fb1 ntpd_intres[30525]: host name not found: 0.debian.pool.ntp.org
Oct 04 18:21:58 fb1 ntpd_intres[30525]: host name not found: 1.debian.pool.ntp.org
Oct 04 18:22:08 fb1 ntpd_intres[30525]: host name not found: 2.debian.pool.ntp.org
Oct 04 18:22:18 fb1 ntpd_intres[30525]: host name not found: 3.debian.pool.ntp.org
- But the ntpq command is giving the saem output as given by ntpq comman in martian fb1 (except for the source servers), that the broadcasting is happening in the same manner:
controls@fb1:~ 0$ ntpq -p
remote refid st t when poll reach delay offset jitter
==============================================================================
192.168.123.255 .BCST. 16 u - 64 0 0.000 0.000 0.000
- On the FE machines side though, the systemd-timesyncd are still unable to read the time signal from fb1 and show the status as idle:
controls@c1bhd:~ 3$ timedatectl
Local time: Mon 2021-10-04 18:34:38 UTC
Universal time: Mon 2021-10-04 18:34:38 UTC
RTC time: Mon 2021-10-04 18:34:38
Time zone: Etc/UTC (UTC, +0000)
NTP enabled: yes
NTP synchronized: no
RTC in local TZ: no
DST active: n/a
controls@c1bhd:~ 0$ systemctl status systemd-timesyncd -l
● systemd-timesyncd.service - Network Time Synchronization
Loaded: loaded (/lib/systemd/system/systemd-timesyncd.service; enabled)
Active: active (running) since Mon 2021-10-04 17:21:29 UTC; 1h 13min ago
Docs: man:systemd-timesyncd.service(8)
Main PID: 244 (systemd-timesyn)
Status: "Idle."
CGroup: /system.slice/systemd-timesyncd.service
└─244 /lib/systemd/systemd-timesyncd
- So the time synchronization is still not working. We expected the FE machined to just synchronize to fb1 even though it doesn't have any upstream ntp server to synchronize to. But that didn't happen.
- I'm (Anchal) working on getting internet access to c1teststand computers.
Digging into mx_stream/daqd_dc errors:
- We went and changed the Restart fileld in /etc/systemd/system/daqd_dc.service on cloned fb1 to 2. This allows the service to fail and stop restarting after two attempts. This allows us to see the real error message instead of the systemd error message that the service is restarting too often. We got following:
controls@fb1:~ 3$ sudo systemctl status daqd_dc -l
● daqd_dc.service - Advanced LIGO RTS daqd data concentrator
Loaded: loaded (/etc/systemd/system/daqd_dc.service; enabled)
Active: failed (Result: exit-code) since Mon 2021-10-04 17:50:25 UTC; 22s ago
Process: 715 ExecStart=/usr/bin/daqd_dc_mx -c /opt/rtcds/caltech/c1/target/daqd/daqdrc.dc (code=exited, status=1/FAILURE)
Main PID: 715 (code=exited, status=1/FAILURE)
Oct 04 17:50:24 fb1 systemd[1]: Started Advanced LIGO RTS daqd data concentrator.
Oct 04 17:50:25 fb1 daqd_dc_mx[715]: [Mon Oct 4 17:50:25 2021] Unable to set to nice = -20 -error Unknown error -1
Oct 04 17:50:25 fb1 daqd_dc_mx[715]: Failed to do mx_get_info: MX not initialized.
Oct 04 17:50:25 fb1 daqd_dc_mx[715]: 263596
Oct 04 17:50:25 fb1 systemd[1]: daqd_dc.service: main process exited, code=exited, status=1/FAILURE
Oct 04 17:50:25 fb1 systemd[1]: Unit daqd_dc.service entered failed state.
- It seemed like the only thing daqd_dc process doesn't like is that mx_stream services are in failed state in teh FE computers. So we did the same process on FE machines to get the real error messages:
controls@fb1:~ 0$ sudo chroot /diskless/root
fb1:/ 0#
fb1:/ 0# sudo nano /etc/systemd/system/mx_stream.service
fb1:/ 0#
fb1:/ 0# exit
- Then I ssh'ed into c1bhd to see the error message on mx_stream service properly.
controls@c1bhd:~ 0$ sudo systemctl daemon-reload
controls@c1bhd:~ 0$ sudo systemctl restart mx_stream
controls@c1bhd:~ 0$ sudo systemctl status mx_stream -l
● mx_stream.service - Advanced LIGO RTS front end mx stream
Loaded: loaded (/etc/systemd/system/mx_stream.service; enabled)
Active: failed (Result: exit-code) since Mon 2021-10-04 17:57:20 UTC; 24s ago
Process: 11832 ExecStart=/etc/mx_stream_exec (code=exited, status=1/FAILURE)
Main PID: 11832 (code=exited, status=1/FAILURE)
Oct 04 17:57:20 c1bhd systemd[1]: Starting Advanced LIGO RTS front end mx stream...
Oct 04 17:57:20 c1bhd systemd[1]: Started Advanced LIGO RTS front end mx stream.
Oct 04 17:57:20 c1bhd mx_stream_exec[11832]: send len = 263596
Oct 04 17:57:20 c1bhd mx_stream_exec[11832]: OMX: Failed to find peer index of board 00:00:00:00:00:00 (Peer Not Found in the Table)
Oct 04 17:57:20 c1bhd mx_stream_exec[11832]: mx_connect failed Nic ID not Found in Peer Table
Oct 04 17:57:20 c1bhd mx_stream_exec[11832]: c1x06_daq mmapped address is 0x7f516a97a000
Oct 04 17:57:20 c1bhd mx_stream_exec[11832]: c1bhd_daq mmapped address is 0x7f516697a000
Oct 04 17:57:20 c1bhd systemd[1]: mx_stream.service: main process exited, code=exited, status=1/FAILURE
Oct 04 17:57:20 c1bhd systemd[1]: Unit mx_stream.service entered failed state.
- c1sus2 shows the same error. I'm not sure I understand these errors at all. But they seem to have nothing to do with timing issues
!
As usual, some help would be helpful |
16371
|
Fri Oct 1 14:25:27 2021 |
yehonathan | Summary | SUS | PRM and BS Angular Actuation transfer function magnitude measurements |
{Paco, Yehonathan, Hang}
We measured the sensing PRMI sensing matrix. Attachment 1 shows the results, the magnitude of the response is not calibrated. The orthogonality between PRCL and MICH is still bad (see previous measurement for reference).
Hang suggested that since MICH actuation with BS and PRM is not trivial (0.5*BS - 0.34*PRM) and since PRCL is so sensitive to PRM movement there might be a leakage to PRCL when we are actuating on MICH. So there may be a room to tune the PRM coefficient in the MICH output matrix.
Attachment 2 shows the sensing matrix after we changed the MICH->PRM coefficient in the OSC output matrix to -0.1.
It seems like it made things a little bit better but not much and also there is a huge uncertainty in the MICH sensing. |
Attachment 1: MICH_PRM_-0.34.png
|
|
Attachment 2: MICH_PRM_-0.1.png
|
|
16370
|
Fri Oct 1 12:12:54 2021 |
Stephen | Update | BHD | ITMY (3002) CAD layout pushed to Box |
Koji requested current state of BHD 3D model. I pushed this to Box after adding the additional SOSs and creating an EASM representation (also posted, Attachment 1). I also post the PDF used to dimension this model (Attachment 2). This process raised some points that I'll jot down here:
1) Because the 40m CAD files are not 100% confirmed to be clean of any student license efforts, we cannot post these files to the PDM Vault or transmit them this way. When working on BHD layout efforts, these assemblies which integrate new design work therefore must be checked for most current revisions of vault-managed files - this Frankenstein approach is not ideal but can be managed for this effort.
2) Because the current files reflect the 40m as built state (as far as I can tell), I shared the files in a zip directory without increasing the revisions. It is unclear whether revision control is adequate to separate [current 40m state as reflected in CAD] from [planned 40m state after BHD upgrade]. Typically a CAD user would trust that we could find the version N assembly referenced in the drawing from year Y, so we wouldn't hesitate to create future design work in a version N+1 assembly file pending a current drawing. However, this form of revision control is not implemented. Perhaps we want to use configurations to separate design states (in other words, create a parallel model of every changed component, without creating paralle files - these configurations can be selected internal to the assembly without a need to replace files)? Or more simply (and perhaps more tenuously), we could snapshot the Box revisions and create a DCC page which notes the point of departure for BHD efforts?
Anyway, the cold hard facts:
- Box location: 40m/40m_cad_models/Solidworks_40m (LINK)
- Filenames: 3002.zip and 3002 20211001 ITMY BHD for Koji presentation images.easm (healthy disregard for concerns about spaces in filenames) |
Attachment 1: 3002_20211001_ITMY_BHD_for_Koji_presentation_images.easm
|
Attachment 2: 40m_upgrade_layout_20200611-ITMY_Beam_Dim.pdf
|
|
16369
|
Thu Sep 30 18:04:31 2021 |
Paco | Summary | Calibration | XARM OLTF (calibration) with three lines |
[anchal, paco]
We repeated the same procedure as before, but with 3 different lines at 55.511, 154.11, and 1071.11 Hz. We overlay the OLTF magnitudes and phases with our latest model (which we have updated with Koji's help) and include the rms uncertainties as errorbars in Attachment #1.
We also plot the noise ASDs of calibrated OLTF magnitudes at the line frequencies in Attachment #2. These curves are created by calculating power spectral density of timeseries of OLTF values at the line frequencies generated by demodulated XARM_IN and ETMX_LSC_OUT signals. We have overlayed the TRX noise spectrum here as an attempt to see if we can budget the noise measured in values of G to the fluctuation in optical gain due to changing power in the arms. We multiplied the the transmission ASD with the value of OLTF at those frequencies as the transfger function from normalized optical gain to the total transfer function value.
It is weird that the fluctuations in transmission power at 1 mHz always crosses the total noise in the OLTF value in all calibration lines. This could be an artificat of our data analysis though.
Even if the contribution of the fluctuating power is correct, there is remaining excess noise in the OLTF to be budgeted. |
Attachment 1: XARM_OLTF_Model_and_Meas.pdf
|
|
Attachment 2: Gmag_ASD_nb_withTRX.pdf
|
|
16368
|
Thu Sep 30 14:13:18 2021 |
Anchal | Update | LSC | HV supply to Xend Green laser injection mirrors M1 and M2 PZT restored |
Late elog, original date Sep 15th
We found that the power switch of HV supply that powers the PZT drivers for M1 and M2 on Xend green laser injection alignment was tripped off. We could not find any log of someone doing it, it is a physical switch. Our only explanation is that this supply might have a solenoid mechansm to shut off during power glitches and it probably did so on Aug 23 (see 40m/16287). We were able to align the green laser using PZT again, however, the maximum power at green transmission from X arm cavity is now about half of what it used to be before the glitch. Maybe the seed laser on the X end died a little. |
16367
|
Thu Sep 30 14:09:37 2021 |
Anchal | Summary | CDS | New way to ssh into c1teststand |
Late elog, original time Wed Sep 29 14:09:59 2021
We opened a new port (22220) in the router to the martian subnetwork which is forwarded to port 22 on c1teststand (192.168.113.245) allowing direct ssh access to c1teststand computer from the outside world using:
Checkout this wiki page for unredadcted info. |
16366
|
Thu Sep 30 11:46:33 2021 |
Ian MacMillan | Summary | Computers | Quantization Noise Calculation Summary |
First and foremost I have the updated bode plot with the mode moved to 10 Hz. See Attachment 1. Note that the comparison measurement is a % difference whereas in the previous bode plot it was just the difference. I also wrapped the phase so that jumps from -180 to 180 are moved down. This eliminates massive jumps in the % difference.
Next, I have two comparison plots: 32 bit and 64bit. As mentioned above I moved the mode to 10 Hz and just excited both systems at 3.4283Hz with an amplitude of 1. As we can see on the plot the two models are practically the same when using 64bits. With the 32bit system, we can see that the noise in the IIR filter is much greater than in the State-space model at frequencies greater than our mode.
Note about windowing and averaging: I used a Hanning window with averaging over 4 neighbor points. I came to this number after looking at the results with less averaging and more averaging. In the code, this can be seen as nperseg=num_samples/4 which is then fed into signal.welch |
Attachment 1: SS-IIR-Bode.pdf
|
|
Attachment 2: PSD_32bit.pdf
|
|
Attachment 3: PSD_64bit.pdf
|
|
16365
|
Wed Sep 29 17:10:09 2021 |
Anchal | Summary | CDS | c1teststand problems summary |
[anchal, ian]
We went and collected some information for the overlords to fix the c1teststand DAQ network issue.
- from c1teststand, c1bhd and c1sus2 computers were not accessible through ssh. (No route to host). So we restarted both the computers (the I/O chassis were ON).
- After the computers restarted, we were able to ssh into c1bhd and c1sus, ad we ran rtcds start c1x06 and rtcds start c1x07.
- The first page in attachment shows the screenshot of GDS_TP screens of the IOP models after this step.
- Then we started teh user models by running rtcds start c1bhd and rtcds start c1su2.
- The second page shows the screenshot of GDS_TP screens. You can notice that DAQ status is red in all the screens and the DC statuses are blank.
- So we checked if daqd_ services are running in the fb computer. They were not. So we started them all by sudo systemctl start daqd_*.
- Third page shows the status of all services after this step. the daqd_dc.service remained at failed state.
- open-mx_stream.service was not even loaded in fb. We started it by running sudo systemctl start open-mx_stream.service.
- The fourth page shows the status of this service. It started without any errors.
- However, when we went to check the status of mx_stream.service in c1bhd and c1sus2, they were not loaded and we we tried to start them, they showed failed state and kept trying to start every 3 seconds without success. (See page 5 and 6).
- Finally, we also took a screenshot of timedatectl command output on the three computers fb, c1bhd, and c1sus2 to show that their times were not synced at all.
- The ntp service is running on fb but it probably does not have access to any of the servers it is following.
- The timesyncd on c1bhd and c1sus2 (FE machines) is also running but showing status 'Idle' which suggested they are unable to find the ntp signal from fb.
- I believe this issue is similar to what jamie ficed in the fb1 on martian network in 40m/16302. Since the fb on c1teststand network was cloned before this fix, it might have this dysfunctional ntp as well.
We would try to get internet access to c1teststand soon. Meanwhile, someone with more experience and knowledge should look into this situation and try to fix it. We need to test the c1teststand within few weeks now. |
Attachment 1: c1teststand_issues_summary.pdf
|
|
16364
|
Wed Sep 29 09:36:26 2021 |
Jordan | Update | SUS | 2" Adapter Ring Parts for SOS Arrived 9/28/21 |
The remaining machined parts for the SOS adapter ring have arrived. I will inspect these today and get them ready for C&B. |
Attachment 1: 20210929_092418.jpg
|
|
16363
|
Tue Sep 28 16:31:52 2021 |
Paco | Summary | Calibration | XARM OLTF (calibration) at 55.511 Hz |
[anchal, paco]
Here is a demonstration of the methods leading to the single (X)arm calibration with its budget uncertainty. The steps towards this measurement are the following:
- We put a single line excitation through the C1:SUS-ETMX_LSC_EXC at 55.511 Hz, amp = 1 counts, gain = 300 (ramptime=10 s).
- With the arm locked, we grab a long timeseries of the C1:LSC-XARM_IN1_DQ (error point) and C1:SUS-ETMX_LSC_OUT_DQ (control point) channels.
- We assume the single arm loop to have the four blocks shown in Attachment #1, A (actuator + sus), plant (mainly the cavity pole), D (detection + electronics), and K (digital control).
- At this point, Anchal made a model of the single arm loop including the appropriate filter coefficients and other parameters. See Attachments #2-3 for the split and total model TFs.
- Our line would actually probe a TF from point b (error point) to point d (control point). We multiplied our measurement with open loop TF from b to d from model to get complete OLTF.
- Our initial estimate from documents and elog made overall loop shape correct but it was off by an overall gain factor. This could be due to wrong assumption on RFPD transimpedance or analog gains of AA or whitening filters. We have corrected for this factor in the RFPD transimpedance, but this needs to be checked (if we really care).
- We demodulate decimated timeseries (final sampling rate ~ 2.048 kHz) and I & Q for both the b and d signals. From this and our model for K, we estimate the OLTF. Attachment #4 shows timeseries for magnitude and phase.
- Finally, we compute the ASD for the OLTF magnitude. We plot it in Attachment #5 together with the ASD of the XARM transmission (C1:LSC-TRX_OUT_DQ) times the OLTF to estimate the optical gain noise ASD (this last step was a quick attempt at budgeting the calibration noise).
- For each ASD we used N = 24 averages, from which we estimate rms (statistical) uncertainties which are depicted by error bands (
) around the lines.
** Note: We ran the same procedure using dtt (diaggui) to validate our estimates at every point, as well as check our SNR in b and d before taking the ~3.5 hours of data. |
Attachment 1: OLTF_Calibration_Scheme.jpg
|
|
Attachment 2: XARM_POX_Lock_Model_TF.pdf
|
|
Attachment 3: XARM_OLTF_Total_Model.pdf
|
|
Attachment 4: XARM_OLTF_55p511_Hz_timeseries.pdf
|
|
Attachment 5: Gmag_55p511_Hz_ASD.pdf
|
|
16362
|
Mon Sep 27 17:04:43 2021 |
rana | Summary | Computers | Quantization Noise Calculation Summary |
I suggest that you
- change the corner frequency to 10 Hz as I suggested last week. This filter, as it is, is going to give you trouble.
- Put in a sine wave at 3.4283 Hz with an amplitude of 1, rather than white noise. In this way, its not necessary to do any subtraction. Just make PSD of the output of each filter.
- Be careful about window length and window function. If you don't do this carefully, your PSD will be polluted by window bleeding.
|
16361
|
Mon Sep 27 16:03:15 2021 |
Ian MacMillan | Summary | Computers | Quantization Noise Calculation Summary |
I have coded up the procedure in the previous post: The result does not look like what I would expect.
As shown in Attachment1 I have the power spectrum of the 32-bit output and the 64-bit output as well as the power spectrum of the two subtracted time series as well as the subtracted power spectra of both. unfortunately, all of them follow the same general shape of the raw output of the filter.
I would not expect quantization noise to follow the shape of the filter. I would instead expect it to be more uniform. If anything I would expect the quantization noise to increase with frequency. If a high-frequency signal is run through a filter that has high quantization noise then it will severely degrade: i.e. higher quantization noise.
This is one reason why I am confused by what I am seeing here. In all cases including feeding the same and different white noise into both filters, I have found that the calculated quantization noise is proportional to the response of the filter. this seems wrong to me so I will continue to play around with it to see if I can gain any intuition about what might be happening. |
Attachment 1: DeltaNoiseSpectrum.pdf
|
|
16360
|
Mon Sep 27 12:12:15 2021 |
Ian MacMillan | Summary | Computers | Quantization Noise Calculation Summary |
I have not been able to figure out a way to make the system that Aaron and I talked about. I'm not even sure it is possible to pull the information out of the information I have in this way. Even the book uses a comparison to a high precision filter as a way to calculate the quantization noise:
"Quantization noise in digital filters can be studied in simulation by comparing the behavior of the actual quantized digital filter with that of a refrence digital filter having the same structure but whose numerical calculations are done extremely accurately."
-Quantization Noise by Bernard Widrow and Istvan Kollar (pg. 416)
Thus I will use a technique closer to that used in Den Martynov's thesis (see appendix B starting on page 171). A summary of my understanding of his method is given here:
A filter is given raw unfiltered gaussian data then it is filtered and the result is the filtered data thus we get the result: where is the raw noise filtered through an ideal filter and is the difference which in this case is the quantization noise. Thus I will input about 100-1000 seconds of the same white noise into a 32-bit and a 64-bit filter. (hopefully, I can increase the more precise one to 128 bit in the future) then I record their outputs and subtract the from each other. this should give us the Quantization error :

and since because they are both running through ideal filters:


and since in this case, we are assuming that the higher bit-rate process is essentially noiseless we get the Quantization noise .
If we make some assumptions, then we can actually calculate a more precise version of the quantization noise:
"Since aLIGO CDS system uses double precision format, quantization noise is extrapolated assuming that it scales with mantissa length"
-Denis Martynov's Thesis (pg. 173)
From this assumption, we can say that the noise difference between the 32-bit and 64-bit filter outputs: is proportional to the difference between their mantissa length. by averaging over many different bit lengths, we can estimate a better quantization noise number.
I am building the code to do this in this file |